OpenStudy (anonymous):

harry has a jar that contains 75% of apple juice and the rest water. The bottle has 1 liter of water. Part A: write an equation in one variable that can be used to find the total number of liters of apple juice and water in the jar. Define the variable used in the equation. Part B:How many liters of apple juice are present in the jar? show your work.

OpenStudy (anonymous):

75% represents unknown quantity of juice. 25% represents one liter of water. can you continue from here?

OpenStudy (anonymous):

yes I think I can

OpenStudy (anonymous):

can you still show me though? just in case

OpenStudy (anonymous):

what do you have so far?

OpenStudy (anonymous):

nothing really.... I am really confused

OpenStudy (anonymous):

if 25% = 1 liter, what does 75% =?

OpenStudy (anonymous):

75% equals the amount of apple juice?

OpenStudy (anonymous):

how much? if 25% is one liter, then 75% is how much apple juice

OpenStudy (anonymous):

how would I figure that out?

OpenStudy (anonymous):



OpenStudy (anonymous):

So 75% equals 3 liters...?

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

so you have determined that you have 1 liter of water and 3 liters of juice.

OpenStudy (anonymous):

so how do I write that in equation form?

OpenStudy (anonymous):

another way of saying liters is to use x as a variable. if x represents 1 liter, how would you represent 3 liters?

OpenStudy (anonymous):

would I use an exponent?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

x =1 2x = 2....

OpenStudy (anonymous):

3x?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

so let's call water W, and juice J. total= xW+3xJ

OpenStudy (anonymous):

that means that the total is equal to 1 part water and three parts juice. You are given water = 1, so plugging in one for x, y=x+3x y=1+3 y=4

OpenStudy (anonymous):

are we done with part a then?

OpenStudy (anonymous):

you've got part A and B there

OpenStudy (anonymous):

oh okay thank you!!